# Bridging Perception and Planning: Towards End-to-End Planning for Signal Temporal Logic Tasks

### v1

@@ I Introduction
Signal Temporal Logic (STL) has emerged as a widely-used specification language in the design of autonomous systems, primarily due to its quantitative robustness semantics and the wealth of temporal operators it offers. It has been successfully applied across a broad range of engineering applications [1, 29]. Particularly, in the field of robotics motion and task planning, STL has garnered significant attention and has been widely used in areas such as aerial robot [25], underwater vehicles [8] and marine surface vessels [7].
@@
Despite its success in robotic task and motion planning, current STL methods remain largely restricted to scenarios where structured system information is available a priori. This structured information typically includes: (i) environmental semantics (e.g., atomic predicates), (ii) the abstracted workspace for discrete-level planning (commonly represented as grids or polytopic regions), and (iii) system dynamics when planning over both tasks and motions. Most existing STL planning approaches, including encoding-based [23, 26], gradient-based [9, 17, 20], and learning-based optimization methods [5, 27], fall into this category, as they all fundamentally rely on predefined structural information at some level.
@@
Current methods typically separate environmental perception, structured information modeling, and STL task planning into distinct stages. Most research focuses on backend planning based on structured information, while the impact of perceiving and abstracting unstructured environmental data is often overlooked. Acquiring structured information usually involves extracting data from the system’s nonlinear dynamics or high-dimensional inputs, with robots often relying on raw sensory data like camera images. Existing methods, however, prevent direct reasoning over such data, making systems brittle when environments change or map annotations are unavailable. An end-to-end paradigm that learns to satisfy STL specifications directly from perceptual inputs could bypass manual discretization, enabling better integration between perception and decision-making. This approach aligns with broader advancements in robotics, where end-to-end learning has proven beneficial.
@@
Building on this perspective, recent studies in autonomous driving [10, 12] and robotic manipulation empowered by vision–language–action (VLA) models [15, 2, 30] show that end-to-end learning paradigms are advantageous in unifying representation and decision-making. Inspired by these findings, we believe that such unification could significantly enhance STL-constrained planning. However, solving STL tasks directly from raw visual observations remains an unexplored challenge. Existing methods still rely on explicit map abstractions or low-dimensional state inputs. This gap motivates our pursuit of a truly end-to-end framework that integrates perception, temporal logic reasoning, and trajectory generation into a single differentiable architecture.
@@
In this work, motivated by recent advancements in autonomous driving, we propose an end-to-end autoregressive transformer framework that ingests synchronized multi-view camera streams along with an STL specification, and outputs a physically feasible trajectory guaranteed to satisfy the prescribed temporal-logic constraints. However, two key challenges impede the realization of such an end-to-end architecture. First is the issue of dataset unavailability. To our knowledge, no public corpus offers synchronized multi-view imagery, formal STL task specifications, and expert trajectories simultaneously, complicating supervised training and objective benchmarking. Second is the complexity of spatio-temporal learning. Specifically, the rich temporal operators and spatial predicates of STL create a highly non-convex, long-horizon optimization landscape, making gradient-based learning more demanding than in conventional trajectory-prediction settings. To address these challenges, our work makes the following contributions:
@@
  • Structure-Aware Mixture of Experts (MoE). We introduce a structure-aware MoE model that decomposes each STL formula into temporally anchored sub-task tokens and routes them to operator-specific experts. This facilitates efficient end-to-end learning of STL-compliant trajectories and enhances logical reasoning by exploiting STL compositionality.
  • Benchmark Dataset and Empirical Validation. We construct a high-fidelity, Gazebo-based benchmark that combines temporally synchronized multi-view camera streams with formally annotated STL specifications and expert trajectories across a range of factory-logistics scenarios. This is the first publicly available corpus designed for supervised, end-to-end STL learning. We demonstrate that the proposed model achieves state-of-the-art STL satisfaction and trajectory feasibility, with improved performance and without incurring additional planning latency compared to a single-expert transformer baseline.
@@ II Related Works
End-to-End Model Paradigm. UniAD [10] pioneered the end-to-end framework in autonomous driving, unifying multiple perception tasks into a single architecture to directly support downstream planning, enhancing planning quality and efficiency. End-to-end designs have since been explored in autonomous driving and robotic manipulation. In autonomous driving, multi-task BEV/Vector transformers and closed-loop differentiable planners unify perception, prediction, and planning [13, 21], while diffusion-based trajectory generators produce multimodal futures and safety-guided sampling [14, 18]. In manipulation, diffusion policies in joint or SE(3) spaces demonstrate strong long-horizon performance [4, 24], and VLA models align semantics with control via instruction grounding and policy distillation [3, 16]. These advancements highlight the advantages of end-to-end approaches, such as better coupling of representation and decision-making, uncertainty-aware generative modeling, and language-conditioned grounding. Building on this, we introduce the first end-to-end solution for solving STL-specified tasks using raw sensory inputs and an STL specification directly.
@@ III Signal Temporal Logic Specifications
As the general STL is highly expressive, its hypothesis class can be overly rich relative to the available data, which makes it difficult for a learning-based model to recover precise semantics and yields unstable gradients during training. To ensure learnability and tractable monitoring, we restrict attention to a widely used finite-horizon fragment that still captures the majority of robotic tasks. Specifically, we adopt a discrete-time semantics with horizon $T\in\mathbb{N}$ and timestamps $t\in\{0,1,\dots,T\}$. All temporal operators are bounded in time, with integer bounds $0\leq a\leq b\leq T$. We further limit temporal nesting depth to $1$: each temporal operator applies directly to a temporal-free Boolean formula. Formally, let atomic predicates be $\mu(x_{t})$ with robustness $\rho_{\mu}[t]\in\mathbb{R}$, and let $\chi$ denote any Boolean combination of atoms using $\wedge,\vee$ (no temporal operators inside). Our fragment $\mathsf{STL}_{\mathrm{B,depth}\leq 1}$ is generated by
@@ IV Our Scenario and Problem Formulation
Furthermore, without loss of generality, we assume that the training and execution environments share the same physical dimensions. However, the positions of task zones and obstacle regions are allowed to vary.
Figure 1: The Overall Architecture of Model
@@ V E2E Structural Moe Model
The overall pipeline of S-MSP is shown in Fig 1, which is a Transformer-based, end-to-end planner that maps multi-view images and a Signal Temporal Logic (STL) task to a future waypoint sequence. The S-MSP model fuses four conditioning pathways: vision, STL text, pose/region context, and a structural encoding of the STL formula and decodes trajectories with a Hierarchical Key-Value (HKV) Mixture-of-Experts (MoE) Transformer. Formally,
$f_{\theta}:\;(\mathcal{I},\,\varphi,\,\mathbf{p}_{0},\,\mathcal{R})\;\mapsto\;\hat{\mathbf{Y}}_{1:T};\hat{\mathbf{y}}_{t}=(x_{t},y_{t},\psi_{t})$ (3)
@@ V-A1 Image Encoder
where d=256d{=}256 is the embedding dimension and $N_{v}$ is the post-aggregation token count.
@@ V-A2 STL Text Encoder
$\mathbf{L}=\mathrm{Proj}(\mathbf{l})\in\mathbb{R}^{B\times m_{\text{text}}\times d},$ (5)
@@ V-A4 Structural STL Encoder
We parse the STL specification $\varphi$ into an operator–predicate AST and linearize it in preorder with a fixed sibling order to obtain a typed token sequence. For each node $m$ we form an embedding by summing token, time-interval, and depth cues:
@@
$\mathbf{s}_{m}$ $=E_{\text{tok}}(\mathrm{id}_{m})\;+\;E_{\text{time}}(\tilde{\Delta}_{m})\;+\;E_{\text{depth}}(d_{m}),$ (7b)
where $\mathrm{id}_{m}$ indexes a unified vocabulary covering operators (F/G/and/or) and targeted region tokens; $\tilde{\Delta}_{m}=[\tilde{a}_{m},\tilde{b}_{m}]\in[0,1]^{2}$ is the normalized time interval (obtained by dividing by the planning horizon), and $d_{m}$ is the AST depth (clamped to $0,D_{\max}{-}1]$). We prepend a learnable $\langle\mathrm{CLS}\rangle$ token and feed $\langle\mathrm{CLS}\rangle;\mathbf{S}]$ to a small Transformer encoder; the $\langle\mathrm{CLS}\rangle$ output serves as the global formula embedding, while per-token outputs are exposed for downstream structural routing.
@@ V-B Autoregressive Trajectory Decoder with STL-Structured HKV–MoE
$\mathbf{M}_{0}\;=\;\big{[}\,\mathbf{V}\;;\;\mathbf{L}\;;\;\mathbf{c_{1},c_{2}}\;\big{]}\in\mathbb{R}^{(N_{v}+1+m_{\text{text}})\times d},$
@@
$\mathbf{h}_{t}\;=\;\mathrm{DecBlock}(\mathbf{q}_{t};\mathbf{M}_{t-1});\hat{\mathbf{y}}_{t}\;=\;W_{\text{head}}\,\mathbf{h}_{t}\in\mathbb{R}^{3}.$
@@
STL-Structured MoE. Beyond the standard Mixture-of-Experts (MoE) architecture, we develop a STL-conditioned expert-routing scheme that harnesses the compositional structure inherent in the temporal logic specification. $\mathbf{S}$ are structural STL tokens (Eq. (7)) denote the per-token structural memory obtained from the structural STL encoder. For each decoding step $t$, we determine the innermost temporal operator whose bounded interval covers $t$, denoted by $\mathrm{op}(t)$, and assign a coarse time band within that interval, $\mathrm{band}(t)\in\{0,\ldots,K{-}1\}$. A deterministic map $b(\cdot)$ combines operator family and time band into one of $B$ routing buckets:
@@ V-C Training Objective
Input : Nominal $\hat{\mathbf{Y}}$; STL formula $\varphi$; horizon $H$; goal-bias $\beta$; attempts $M$; limits $(d_{\max},\Delta\theta_{\max})$
Output : Repaired $\tilde{\mathbf{Y}}$
1 $\tilde{\mathbf{Y}}\leftarrow\hat{\mathbf{Y}}$; $t\leftarrow 1$
@@
3 $\rho\leftarrow\texttt{robustness}(\tilde{\mathbf{Y}}_{t:t+H},\,\varphi)$
[…]
5    $\tau\leftarrow t$; initialize tree $\mathcal{T}$ with root $z_{\tau-1}$
@@
7         Sample target $q$: with prob. $\beta$ draw near $\{\hat{p}_{\tau},\dots,\hat{p}_{\tau+H}\}$ or predicate-relevant sites of $\varphi$; else uniform
@@ 8       $z_{\text{near}}\leftarrow\texttt{nearest}(\mathcal{T},q)$
9       $z_{\text{new}}\leftarrow\texttt{steer}(z_{\text{near}},q;d_{\max},\Delta\theta_{\max})$
[…]
11            insert $z_{\text{new}}$ and edge into $\mathcal{T}$
12          if $z_{\text{new}}$ reaches time $\tau{+}H$ then
13             $\rho_{\text{seg}}\leftarrow\texttt{robustness}(\texttt{path}_{\tau:\tau+H},\varphi)$
14               evaluate $\mathcal{J}$ via (16) and update the incumbent best
15
@@ 20 $t\leftarrow t+1$
21
Return repaired $\tilde{\mathbf{Y}}$
@@ V-D RRT-Guided Local Replanning with STL Awareness
Given a nominal trajectory $\hat{\mathbf{Y}}=\{(x_{t},y_{t},\theta_{t})\}_{t=1}^{T}$, we run a rollout-and-check loop. If a forward lookahead at time $\tau$ indicates (i) collision/kinematic violation, or (ii) STL robustness of the local window is negative, we trigger short-horizon RRT-style replanning on $\tau,\tau{+}H]$. Expansion is goal-biased: with probability $\beta$ targets are sampled near nominal anchors $\{\hat{p}_{t}\}_{t=\tau}^{\tau+H}$, otherwise uniformly. Candidate edges must satisfy local limits $\|p_{t}-p_{t-1}\|\!\leq\!d_{\max}$ and $|\theta_{t}-\theta_{t-1}|\!\leq\!\Delta\theta_{\max}$. Among feasible branches that reach $\tau{+}H$, we select the path minimizing
$\mathcal{J}=\sum_{t=\tau}^{\tau+H}\!\Big{(}\lambda_{1}\|p_{t}-\hat{p}_{t}\|_{2}^{2}+\lambda_{2}\,\kappa_{t}^{2}\Big{)}\;-\;\lambda_{3}\,\tilde{\rho}_{\varphi}\!\big{(}\mathbf{Y}_{\tau:\tau+H}\big{)},$ (16)
where $p_{t}=(x_{t},y_{t})$, $\kappa_{t}$ is a discrete curvature proxy, and $\tilde{\rho}_{\varphi}$ is a (smoothed) STL robustness of the segment (higher is better). The repaired segment is optionally smoothed (e.g., cubic spline) and spliced back into $\hat{\mathbf{Y}}$. The full procedures, termed TSP, is shown in Algorithm 1.
label single-F single-G only-AND only-OR only(AND–OR) All (ID) OOD-1 (All) OOD-2 (All) OOD-3 (All)

### v2

@@ I Introduction
Signal Temporal Logic (STL) has emerged as a widely-used specification language in the design of autonomous systems, primarily due to its quantitative robustness semantics and the wealth of temporal operators it offers. It has been successfully applied across a broad range of engineering applications [1, 31, 34, 33]. Particularly, in the field of robotics motion and task planning, STL has garnered significant attention [8, 32, 26, 7].
@@
Despite its success in robotic task and motion planning, current STL methods remain largely restricted to scenarios where structured system information is available a priori. This structured information typically includes: (i) environmental semantics (e.g., atomic predicates), (ii) the abstracted workspace for discrete-level planning (commonly represented as grids or polytopic regions), and (iii) system dynamics when planning over both tasks and motions. Most existing STL planning approaches, including encoding-based [24, 25], gradient-based [9, 17, 21] and learning-based optimization methods [5, 27, 28, 10, 19], fall into this category, as they all fundamentally rely on predefined structural information at some level.
@@
Current methods typically separate environmental perception, structured information modeling, and STL task planning into distinct stages. Consequently, most efforts emphasize planning over pre-structured abstractions, while the influence of perceptual uncertainty and abstraction quality from unstructured observations is comparatively underexplored. In practice, structured information must be inferred from nonlinear system dynamics and high-dimensional sensory inputs (e.g., images), yet conventional pipelines do not support direct reasoning over such raw data, leading to brittle performance under distribution shift or missing map annotations. An end-to-end framework that learns to satisfy STL specifications directly from perceptual inputs can reduce manual discretization and better couple perception with decision-making, consistent with recent successes of end-to-end learning in robotics.
@@
Building on this perspective, recent studies in autonomous driving [11, 13] and robotic manipulation empowered by vision–language–action (VLA) models [2, 35] show that end-to-end learning paradigms are advantageous in unifying representation and decision-making. Inspired by these findings, we believe that such unification could significantly enhance STL-constrained planning. However, solving STL tasks directly from raw visual observations remains an unexplored challenge. Existing methods still rely on explicit map abstractions or low-dimensional state inputs. This gap motivates our pursuit of a truly end-to-end framework that integrates perception, temporal logic reasoning, and trajectory generation into a single differentiable architecture.
@@
In this work, we propose an end-to-end autoregressive transformer framework that ingests synchronized multi-view camera streams along with an STL specification, and outputs a physically feasible trajectory guaranteed to satisfy the prescribed temporal-logic constraints. However, two key challenges impede the realization of such an end-to-end architecture. First is the issue of dataset unavailability. To our knowledge, no public corpus offers synchronized multi-view imagery, formal STL task specifications, and expert trajectories simultaneously, complicating supervised training and objective benchmarking. Second is the complexity of spatio-temporal learning. Specifically, the rich temporal operators and spatial predicates of STL create a highly non-convex, long-horizon optimization landscape, making gradient-based learning more demanding than in conventional trajectory-prediction settings. To address these challenges, our work makes the following contributions:
@@
  • Structure-Aware Mixture of Experts. We introduce a structure-aware MoE model that decomposes each STL formula into temporally anchored sub-task tokens and routes them to operator-specific experts. This facilitates efficient end-to-end learning of STL-compliant trajectories and enhances logical reasoning by exploiting STL compositionality.
  • Benchmark Dataset and Empirical Validation. We construct a high-fidelity, Gazebo-based benchmark that combines temporally synchronized multi-view camera streams with formally annotated STL specifications and expert trajectories across a range of factory-logistics scenarios. This is the first publicly available corpus designed for supervised, end-to-end STL learning. We then demonstrate that the proposed model achieves state-of-the-art STL satisfaction and trajectory feasibility, with improved performance and without incurring additional planning latency compared to a single-expert transformer baseline.
@@ II Related Works
End-to-End Model Paradigm. UniAD [11] pioneered the end-to-end framework in autonomous driving, unifying multiple perception tasks into a single architecture to directly support downstream planning, enhancing planning quality and efficiency. End-to-end designs have since been explored in autonomous driving and robotic manipulation. In autonomous driving, multi-task BEV/Vector transformers and closed-loop differentiable planners unify perception, prediction, and planning [14, 22, 30], while diffusion-based trajectory generators produce multimodal futures and safety-guided sampling [15, 18]. In manipulation, diffusion policies in joint or SE(3) spaces demonstrate strong long-horizon performance [4], and VLA models align semantics with control via instruction grounding and policy distillation [3, 16]. These advancements highlight the advantages of end-to-end approaches, such as better coupling of representation and decision-making, uncertainty-aware generative modeling, and language-conditioned grounding. Building on this, we introduce the first end-to-end solution for solving STL-specified tasks using raw sensory inputs and an STL specification directly.
@@ III Signal Temporal Logic Specifications
As the general STL is highly expressive, its hypothesis class can be overly rich relative to the available data, which makes it difficult for a learning-based end-to-end model to recover precise semantics and yields unstable gradients during training. To ensure learnability and tractable monitoring, we restrict attention to a widely used finite-horizon fragment that still captures the majority of robotic tasks. Specifically, we adopt a discrete-time semantics with horizon $T\in\mathbb{N}$ and timestamps $t\in\{0,1,\dots,T\}$. All temporal operators are bounded in time, with integer bounds $0\leq a\leq b\leq T$. We further limit temporal nesting depth to $1$: each temporal operator applies directly to a temporal-free Boolean formula. Formally, let atomic predicates be $\mu(x_{t})$ with robustness $\rho_{\mu}[t]\in\mathbb{R}$, and let $\chi$ denote any Boolean combination of atoms using $\wedge,\vee$ (no temporal operators inside). Our fragment $\mathsf{STL}_{\mathrm{B,depth}\leq 1}$ is generated by
@@ IV Our Scenario and Problem Formulation
  • Labeled target areas: To enable the model to learn the correspondence between target regions and STL task specifications from data, we mark different regions with distinct colors. Note that the model input consists only of multi-view concentrated image observations, without any explicit geometric annotations (e.g., region coordinates or masks).
@@
Furthermore, without loss of generality, we assume that the training and execution environments share the same physical dimensions. However, the positions of task zones and obstacle regions are allowed to vary, thereby generating diverse scenarios.
Figure 1: The Overall Architecture of Model.
@@ V E2E Structural Moe Model
The overall pipeline of S-MSP is shown in Fig 1, which is a Transformer-based, end-to-end planner that maps multi bird-eye-view images and a Signal Temporal Logic (STL) task to a future waypoint sequence. The S-MSP model fuses four conditioning pathways: vision, STL text, pose/region context, and a structural encoding of the STL formula and decodes trajectories with a Hierarchical Key-Value (HKV) Mixture-of-Experts (MoE) Transformer. Formally,
$f_{\theta}:\;\mathcal{I}\times\varphi\times\mathbf{p}_{0}\times\mathcal{R}\;\mapsto\;\hat{\mathbf{Y}}_{1:T};\hat{\mathbf{y}}_{t}=(x_{t},y_{t},\psi_{t})$ (3)
@@ V-A1 Image Encoder
where $d$ is the embedding dimension and $N_{v}$ is the post-aggregation token count.
@@ V-A2 STL Text Encoder
$\mathbf{L}=\mathrm{Proj}(\mathbf{l})\in\mathbb{R}^{m_{\text{text}}\times d},$ (5)
@@ V-A4 Structural STL Encoder
We parse the STL specification $\varphi$ into an operator–predicate abstract syntax tree(AST) and linearize it in preorder with a fixed sibling order to obtain a typed token sequence. For each node $m$ we form an embedding by summing token, time-interval, and depth cues:
@@
$\mathbf{s}_{m}$ $=E_{\text{tok}}(\mathrm{id}_{m})\;+\;E_{\text{interval}}(\tilde{\Delta}_{m})\;+\;E_{\text{depth}}(d_{m}),$ (7b)
where $L$ denotes the number of subformulaes, $\mathrm{id}_{m}$ indexes a unified vocabulary covering operators (F/G/and/or) and targeted region tokens; $\tilde{\Delta}_{m}=[\tilde{a}_{m},\tilde{b}_{m}]\in[0,1]^{2}$ is the normalized time interval (obtained by dividing by the planning horizon), and $d_{m}$ is the AST depth. We prepend a learnable $\langle\mathrm{CLS}\rangle$ token and feed $\langle\mathrm{CLS}\rangle;\mathbf{S}]$ to a small Transformer encoder; the $\langle\mathrm{CLS}\rangle$ output serves as the global formula embedding, while per-token outputs are exposed for downstream structural routing.
@@ V-B Autoregressive Trajectory Decoder with STL-Structured HKV–MoE
$\mathbf{M}_{0}\;=\;\big[\,\mathbf{V}\;;\;\mathbf{L}\;;\;\mathbf{c_{1},c_{2}}\;\big]\in\mathbb{R}^{(N_{v}+m_{\text{text}}+2)\times d},$
@@
$\mathbf{h}_{t}\;=\;\mathrm{DecBlock}(\mathbf{q}_{t};\mathbf{M}_{t-1});\quad\hat{\mathbf{y}}_{t}\;=\;W_{\text{head}}\,\mathbf{h}_{t}\in\mathbb{R}^{3}.$
@@
STL-Structured MoE. Beyond the standard MoE architecture, we develop a STL-conditioned expert-routing scheme that harnesses the compositional structure inherent in the temporal logic specification. $\mathbf{S}$ are structural STL tokens (Eq. (7)) denote the per-token structural memory obtained from the structural STL encoder. For each decoding step $t$, we determine the innermost temporal operator whose bounded interval covers $t$, denoted by $\mathrm{op}(t)$, and assign a coarse time band within that interval, $\mathrm{band}(t)\in\{0,\ldots,K{-}1\}$, where $K$ is the pre-defined interval gap. A deterministic map $b(\cdot)$ combines operator family and time band into one of $B$ routing buckets:
@@ V-C Training Objective
Input : Nominal trajectory $\hat{\mathbf{Y}}=\{\hat{y}_{1},\dots,\hat{y}_{T}\}$; STL formula $\varphi$; horizon $H$; goal-bias $\beta$; sampling times $M$; limits $(d_{\max},\Delta\theta_{\max})$
Output : Repaired trajectory $\widetilde{\mathbf{Y}}$
1
[…]
@@
4 $H_{t}=\min(H,\,T-t-1)$
5 if $CK(\widetilde{\mathbf{Y}},t)$ or $\rho(\widetilde{\mathbf{Y}}_{t:t+H_{t}},\,\varphi)<0$ then
6      Initialize tree $\mathcal{T}$ with root $z_{0}=\widetilde{y}_{t-1}$; set $z_{H_{t}+1}=\widetilde{y}_{t+H_{t}+1}$
7
@@ 4 $H_{t}=\min(H,\,T-t-1)$
9         Sample target $q$: with prob $\beta$ sample near anchors $\{\hat{p}_{\tau},\dots,\hat{p}_{\tau+H_{t}}\}$; else sample uniformly
@@ 10       $z_{\text{near}}\leftarrow\texttt{nearest}(\mathcal{T},q)$
11       $z_{\text{new}}\leftarrow\texttt{run}(z_{\text{near}},q;d_{\max},\Delta\theta_{\max})$
12
13        insert $z_{\text{new}}$ and edge $(z_{\text{near}},z_{\text{new}})$ into $\mathcal{T}$
14
15    $\mathbf{Y}_{new}=\texttt{path}(\mathcal{T},z),$ with min $\mathcal{J}$ via (16), $\forall z$ satisfies $\texttt{height(z)}=H_{t}$ and $feasible(z,z_{H_{t}+1})$
@@ 17 $t\leftarrow t+1$
18
19return $\widetilde{\mathbf{Y}}$
@@ V-D RRT-Guided Local Replanning with STL Awareness
Given a nominal trajectory $\hat{\mathbf{Y}}=\{(x_{t},y_{t},\theta_{t})\}_{t=1}^{T}$, we run a rollout-and-check loop. If a forward lookahead at time $\tau$ indicates (i) collision/kinematic violation, denoted by $CK(Y,\tau)$, or (ii) STL robustness of the local window is negative, we trigger short-horizon RRT-style replanning on $\tau,\tau{+}H]$. Expansion is goal-biased: with probability $\beta$ targets are sampled near nominal anchors $\{\hat{p}_{t}\}_{t=\tau}^{\tau+H}$, otherwise uniformly. Candidate edges must satisfy local limits $\|p_{t}-p_{t-1}\|\!\leq\!d_{\max}$ and $|\theta_{t}-\theta_{t-1}|\!\leq\!\Delta\theta_{\max}$. Among feasible branches that reach $\tau{+}H$, we select the path minimizing
$\mathcal{J}=\sum_{t=\tau}^{\tau+H}\!\Big(\lambda_{1}\|p_{t}-\hat{p}_{t}\|_{2}^{2}+\lambda_{2}\,\kappa_{t}^{2}\Big)\;-\;\lambda_{3}\,\tilde{\rho}_{\varphi}\!\big(\widetilde{\mathbf{Y}}_{\tau:\tau+H}\big),$ (16)
where $p_{t}=(x_{t},y_{t})$, $\kappa_{t}$ is a discrete curvature proxy, and $\tilde{\rho}_{\varphi}$ is the STL robustness of the segment (higher is better). The repaired segment is optionally smoothed (e.g., cubic spline) and spliced back into $\hat{\mathbf{Y}}$. The full procedures, termed TSP, is shown in Algorithm 1.
Method single-F single-G only-AND only-OR only(AND–OR) All (ID) OOD-1 (All) OOD-2 (All) OOD-3 (All)